# Diffusion MRI with Machine Learning

### v1

@@ I-C Why machine learning may help
Because of its unique advantages such as those mentioned above, machine learning may be able to improve the accuracy, robustness, generalizability, and computational speed for many dMRI analysis tasks. In the following sections, we describe some of the recent studies that have investigated this potential.
II Prior works on dMRI analysis with machine learning
II-A Micro-structure mapping
@@
There have been great efforts and much progress in developing bio-physical models that relate the measured dMRI signal to the micro-structure of brain tissue [14, 13]. Advanced models depend on specialized measurement protocols and rely on computationally complex numerical optimization methods. The estimation problem to be solved is typically nonlinear, sensitive to measurement noise and initialization, and may be unstable and computationally expensive [47, 48]. Accuracy and precision of model fitting is hard to verify and can significantly depend on the optimization algorithm used to perform the fitting [47]. To avoid local minima, some works have resorted to computationally expensive approaches such as grid search, multi-start methods, cascade optimization, and stochastic optimization techniques [49, 50, 48, 47]. There has been a great interest in developing methods, such as those based on dictionary matching [51, 52], to reduce the computational time without compromising the estimation accuracy. A good example of such methods is AMICO [52], which reformulates the micro-structure estimation equations as convex optimization problems that can be solved much faster.
@@
In recent years, machine learning has increasingly been applied to these estimation tasks. Overall, five main justifications have been cited for preferring a machine learning-based method for this application: (1) Unlike the conventional methods that presume a known fixed relationship between the dMRI measurements and the target parameter, machine learning methods can learn this relationship from data [21, 26]. For certain micro-structural parameters, such as residence time of water inside axons, it is believed that existing mathematical models that express the relationship with the dMRI signal are either too simplistic or intractable, and that the existing numerical forward models are computationally too expensive to be used in estimating the parameters from data [43, 25]. Fick et al. show that, for axonal diameter estimation, existing signal models fail outside a narrow range of diameters while a machine learning method can achieve accurate estimation for the whole range of diameters in the data [21]. (2) Conventional methods often involve several steps, where there is no feedback from the later steps to the earlier steps [26, 47]. As a result, these methods are hard to design and optimize. Machine learning methods such as deep neural networks, on the other hand, may be optimized as a single processing step. (3) Machine learning methods can be much faster than numerical optimization routines. Again, such is the case for neural networks that run on GPUs [34, 35, 36]. (4) Standard methods perform the model fitting in a voxel-wise manner, which fails to exploit the spatial correlations to improve the estimation accuracy. Many machine learning models can effectively learn complex spatial correlations directly from data [53]. (5) With machine learning methods it is typically much easier to incorporate prior knowledge such as constraints on the parameters to be estimated or additional MRI contrasts.
Tables I, II, and III list some of the recent works on this topic. We summarize our main observations and conclusions from our study of these methods below.
@@
TABLE I: A summary of recent machine learning methods for estimating the diffusion tensor or its derived parameters such as FA and MD.
 Reference
Summary
 [76]
This work reported higher reconstruction accuracy and precision than standard estimation methods from as few as 3 diffusion-weighted measurements. With a model trained on healthy subjects, FA-based delineation of brain tumors was more accurate than with a standard method.
[28]
This method computes the residuals between under-sampled ($n=6$) data and high-quality targets computed from densely-sampled data. The diffusion tensor is then computed with a standard method. Compared with conventional estimation methods, it reduced the number of measurements by a factor of 3.3-4.6. It improved DTI estimation, DTI-based tractography, and tract-specific analysis.
[82]
This work proposes a technique to handle the measurements acquired with different schemes. It is simply based on projecting the q-space data into orthogonal planes.
[38]
@@
This work reports accurate estimation of FA, MD, and the main diffusion tensor eigenvector from as few as six measurements. The model trained on healthy HCP brains works well on pathological brains. Evaluations included qualitative and quantitative tractography assessment.
[74]
@@
This work has reported reconstruction of fetal DTI with unprecedented accuracy. The method was trained using synthetic data that were generated with a novel pipeline using both fetal in-utero data as well as scans of premature neonates. Evaluations included quantitative comparisons with conventional methods as well as detailed assessment by three medical experts.
[60]
This study has achieved accurate cardiac DTI estimation from six dMRI measurements. It includes a U-Net that estimates the data for additional directions that the six measured. The method is evaluated on ex-vivo data.
[32]
The model (a residual CNN) trained with purely synthetic data performs well on synthetic as well as in-vivo test data for cardiac DTI estimation. It has superior accuracy and precision compared with standard estimation methods and is less prone to predicting implausible values.
[90]
Deep learning methods can reconstruct cardiac DTI with k-space under-sampling rates of up to 4 without any significant quality reduction compared to reference. A transformer network is shown to achieve superior results than a CNN.
[91]
@@
The proposed method, a patch-wise CNN, reduces the required number of measurements by a factor of two compared with standard estimation methods. The estimated DTI maps were used to trace major white matter tracts with high accuracy.
[33]
@@
This work showed that transformer networks can learn the spatial correlations in dMRI signal and in diffusion tensor. They can reconstruct the diffusion tensor with superior accuracy while reducing the required number of measurements by a factor of 5-15. Evaluations include tractography and structural connectivity.
[71]
Various deep neural networks are evaluated and shown to achieve substantially lower reconstruction error for DTI indices such as FA, MD, and radial diffusivity.

### v2

@@ IV Prior works on dMRI analysis with machine learning
Due to its unique advantages mentioned above, machine learning has the potential to improve the accuracy, robustness, generalizability, and computational speed for many dMRI analysis tasks. This section describes some of the recent studies that have investigated this potential. The diagram in Figure 1 shows the outline of the methods covered in this section.
Figure 1: Outline of the 10 classes of methods that have been surveyed in Section IV.
IV-A Data preprocessing
Diffusion-weighted MRI measurements suffer from artifacts and imperfections that can influence the results of downstream analysis steps and the conclusions drawn from the data. Therefore, proper data preprocessing is an essential first step for dMRI analysis [9]. This section describes the recent progress in using machine learning methods to improve pre-processing of dMRI data.
IV-A1 Denoising
Measurement noise in dMRI has a spatially-varying distribution and depends on factors such as acquisition and reconstruction methods. It is typically approximated using Rician or non–central chi–square distributions [64, 65]. The noise is stronger at high b-values that are essential for estimating fiber orientation and multi-compartment models, and it is of high practical importance because it can influence the analysis results in subtle but significant and complex ways [66, 26, 27]. Denoising methods based on the principal components analysis and non-local means have been popular in dMRI [67, 68]. These methods were often adapted to the dMRI setting in innovative ways. For example St-Jean et al. [69] combined the non-local matching with dictionary learning. In order to adapt the non-local means method to curved non-Euclidean domain of dMRI signals, Chen et al. [70] developed a method based on convolutions on graphs that enabled computing multi-scale features from dMRI data.
In recent years, several neural network-based image denoising algorithms ave been proposed by the computer vision society. These methods, such as Noise2Noise [71], Noise2Void [72], and Noise2Self [73] have often surpassed the performance of classical denoising algorithms. They offer several advantages that are especially useful for medical imaging applications including dMRI: (1) They do not need to model the image prior; (2) They do not need to know the noise distribution; (3) They do not require clean images or paired noisy-clean images for training. The ideas presented in these methods have been adopted to develop dMRI denoising techniques in a few recent studies [74, 75]. Fadnavis et al. proposed Patch2Self, which is based on learning locally-linear relations between different dMRI volumes [74]. Patch2Self predicts the signal in one of the q-space elements (i.e., one channel in a 4D representation of the dMRI data) based on the other elements. In order to exploit spatial correlations, the model uses 3D spatial patches as input and target. Effectiveness of Patch2Self was demonstrated in a study on spinal cord [76], where Patch2Self improved FA repeatability and multiple sclerosis (MS) lesion conspicuity. SDnDTI [75] is another self-supervised method that uses a CNN model. It is specially tailored to single-shell data for DTI estimation. Given a set of N dMRI volumes (each acquired with a different gradient direction but the same gradient strength), SDnDTI selects subsets of six volumes and denoises them. This is done by training the CNN to take these six volumes as input and predict a low-noise target that is computed from all N volumes. This is repeated until all N volumes are denoised. Experiments show that SDnDTI is superior to standard methods such as those based on PCA and non-local means.
IV-A2 Artifact correction
Diffusion MRI measurements can also suffer from various artifacts such as those caused by off-resonance fields, subject motion, and Gibbs ringing [25, 77]. In the past two decades, much research has been devoted to computational methods for retrospective correction of these artifacts. A recent review of these methods has been written by Tax et al. [9]. Machine learning methods have also been proposed to address these artifacts.
Much of the distortion correction algorithms for dMRI, such as those targeting susceptibility-induced distortions, are based on image registration. Image registration for dMRI data is especially challenging because of several factors: (1) Low SNR, especially at higher b-values; (2) Different appearance of dMRI volumes of the same brain acquired with different gradient strengths and directions, which is again exacerbated for higher b values; (3) Difficulty of finding a reference image for registration since all dMRI volumes are affected by distortions; (4) Complex interaction of various artifacts such as signal loss that can further challenge the registration process. Standard methods for distortion correction, such as topup [78], use iterative optimization to compute the displacement field. Deep learning-based registration methods have the potential to address these challenges because they can be trained in an unsupervised manner on distorted data.
Unsupervised [79, 80, 81], semi-supervised [82], and supervised [19] deep learning-based registration methods have been applied on reversed phase-encoding dMRI data to estimate a displacement field for distortion correction. One study [82] combined unsupervised training with supervised training using the displacement fields computed by topup as the target. Experiments showed that this semi-supervised method led to better results than purely supervised and purely unsupervised techniques. Most prior works have shown that deep learning techniques match the accuracy of the conventional methods such as topup while being orders of magnitude faster [79, 81, 82, 80, 83]. One study used distortion-free volumes obtained with point-spread-function encoded EPI as registration target and achieved more accurate distortion correction than topup [19]. The input to the network, which was a 2D Fully Convolutional Network (FCN), consisted of distorted dMRI volumes as well as an anatomical T2-weighted image, and the network output was undistorted dMRI volumes. Machine learning methods also have good generalizability. Studies have shown that deep learning models trained on normal brains work well on abnormal brains [19, 80] and on different (left-right) phase encoding directions [80].
Most deep learning-based methods have used the dMRI volumes as input for computing the displacement fields for distortion correction. Qiao et al., on the other hand, propose to use fiber orientation distribution (FOD) as input [84, 85]. They argue that the information in the individual dMRI data volumes is not sufficient to resolve detailed white matter structures that are essential for distortion correction in areas such as the brainstem [84]. Instead, they compute FOD maps from dMRI volumes acquired with opposite phase encodings and train an FCN to compute the displacement fields. This method achieved better results than topup and two alternative deep learning methods in brainstem and white matter. Another interesting work is that of Schilling et al. [83, 86], which enables susceptibility-induced distortion correction when scans with reversed phase-encodings are not available. They have developed deep learning models to generate an undistorted b0 (non-diffusion-weighted) volume given (i) a distorted b0 volume and (ii) an anatomical (T1-weighted) image. The undistorted and distorted images can then be used to compute the displacement field with topup. Their methods are based on FCNs and generative adversarial networks (GANs, [87]).
Deep learning methods have been developed to address other types of artifacts in dMRI. Ayub et al. have proposed a method based on variational auto-encoders for in-painting cropped dMRI volumes [88]. The model, trained on artificially-cropped volumes and tested on real-world cropped data, was able to correctly inpaint the cropped data. Deep learning methods have been successfully applied to remove Gibbs artifacts in dMRI data. One study trained a CNN on T2-weighted images to estimate Gibbs ringing artifacts [89]. The trained model removed Gibbs artifacts in dMRI volumes as well. Another work trained a CNN on natural images with simulated artifacts and showed that the trained model effectively suppressed Gibbs artifacts in dMRI data and in derived parameter maps such as fractional anisotropy (FA) images [90].
IV-A3 Automatic quality control
Quality control in dMRI data processing pipelines can be challenging because of the diversity of artifacts and large data sizes. A number of deep learning methods have been proposed for automatic quality control in large dMRI studies. Ahmad et al. [91] trained a CNN to classify dMRI volumes as normal or containing artifacts. The artifacts considered in their study included motion-induced signal dropout, inter-slice instability, ghosting, chemical shift and susceptibility-induced distortions. Their method had an accuracy of 92% on a pool of seven datasets with different acquisition protocols. Another work used a mix of real and simulated data to train a quality control method, based on CNNs, for detecting intra-volume motion and signal dropout [92]. To detect motion-corrupted volumes in neonatal dMRI volumes, Kelly et al. trained a random forest on the outputs of an ensemble of CNNs and achieved classification accuracies of well above 90% [93]. Samani et al. also trained a CNN to detect a range of artifacts such as motion artifacts, ghosting, susceptibility-induced artifacts, and chemical shift [94]. They reported a detection accuracy of 98% on multi-scanner dMRI data. Both [92] and [94] used transfer learning to train their networks.
IV-B Quality enhancement, super-resolution, and data interpolation
Increasing the spatial resolution or q-space sampling density during image acquisition would require increasing the scan time. As an alternative, recent studies have proposed to use computational methods. Much of the proposed methodology have been inspired by deep learning-based super-resolution techniques. However, often these standard deep learning techniques have to be significantly modified to handle the dMRI data. A summary of these works is provided below.
IV-B1 Interpolation in q space (i.e., angular space)
Several works have used neural networks to enhance the angular resolution of q-space data [95, 96, 97, 98]. Yin et al. used a method based on sparse representation of the dMRI signal and a neural network built on 1D convolutions to map low angular resolution data to high angular resolution data. This method improved the FOD reconstruction accuracy for complex fiber configurations [96]. In another study, Ren et al. have developed a method for predicting dMRI data for arbitrary gradient directions [99]. The input to this model includes only T1-weighted, T2-weighted, and b0 (non-diffusion-weighted) images. Methodologically, the model is a GAN, where both the generator and the discriminator are conditioned on the gradient direction and strength. Experiments are performed with training and test data from the Human Connectome Project (HCP). The results show that the synthesized data can be used to estimate DKI, Neurite Orientation Dispersion and Density Imaging model (NODDI, [100]), and FOD.
Lyon et al. proposed a parametric continuous convolutional network for angular super-resolution of dMRI data [98]. Their method improved the results of fixel-based analysis and estimation of NODDI. Another work has proposed a flexible model, based on a recurrent convolutional neural network (RCNN), that can predict the unobserved data for arbitrary target b-vectors [101]. To achieve this flexibility, the model follows an encoder-decoder design where the encoder uses the gradient table information of the measured data as additional input. The decoder, then, takes the latent representation generated by the encoder and the gradient table of the target (unobserved) data to predict the unobserved data. This model accurately upsampled the dMRI data by a factor of 8, and it was more accurate than predictions based on spherical harmonic interpolation and a 1D variant of the model architecture. At higher sub-sampling rates, spatial information study improved the accuracy of angular interpolation [101]. A more recent study proposed a geometric deep learning method to predict unobserved q-space data for arbitrary target acquisition schemes such as single/multi-shell and grid-based schemes [102]. Experiments showed that this method was superior to model-based (i.e., non-machine learning) techniques and it improved the estimation accuracy for tissue microstructure and FOD.
IV-B2 Interpolation in voxel space
Different from the methods above, a number of works have attempted to increase the spatial resolution of the dMRI data in voxel space [103, 104]. It is also possible to jointly improve the spatial resolution and q-space resolution, as proposed in super-resolved q-space deep learning (SR-q-DL) framework [105]. In this framework, patches from the source low-resolution and the target high-resolution dMRI volumes are represented in dictionaries and a neural network is trained to map the representation of the source signal to that of the target signal. This method significantly improved the accuracy of estimating the NODDI parameters from low-quality dMRI data. In a more recent work, Spears and Fletcher [106] have argued that prior voxel-space super-resolution methods such as [47] are not ideal for tractography, where a continuous FOD field is desired. Instead, they have proposed a method based on continuous and differentiable signal representations with neural networks. They show that their new method achieves state of the art FOD estimation and tractography on HCP data.
IV-B3 Enhancement of derivative maps
A number of works have proposed to enhance the quality/resolution of the derived parameter maps. Zeng et al. developed a method to enhance the resolution of FOD maps [107]. Their method used the FOD estimated from low angular resolution single-shell data to predict the FOD estimated from multi-shell high angular resolution data. Experiments on clinical quality data showed that this method significantly improved the tractography and structural connectivity assessment. A similar study has been reported by Lucena et al., [108], where the authors trained CNNs to map the FODs computed from single-shell data to those computed from multi-shell data. The method worked well and showed good generalizability to different scanners and data acquisition protocols. Ye et al. [109] proposed to combine a learning-based parameter mapping method similar to [110] with a super-resolution method similar to [111] in a joint framework to estimate high-resolution NODDI maps from dMRI measurements with low resolution in both voxel space and q-space.
IV-B4 Image quality transfer
In a series of works, Alexander et al. have developed and promoted the notion of image quality transfer (IQT), which uses machine learning methods such as random forests and neural networks to improve the quality of dMRI parameter maps [112, 113]. The goal of IQT is to use high-quality data (such as the HCP data) to improve the quality of regular dMRI data that are typically obtained in clinical applications. The method works by learning a regression function to map the low-quality dMRI patches to high-quality patches. This idea has been successfully applied to increase the spatial resolution of diffusion tensor images and to estimate NODDI parameters from low-quality single-shell measurements. Furthermore, it has been shown that on low-quality data, IQT enables tractography-based reconstruction of thin fiber tracts that can typically only be reconstructed using specialized data.
IV-B5 Estimation/imputation of unobserved or missing data
Estimating unobserved q-space shells from observed ones has been reported by several works [114, 115, 95]. Murray et al. used neural networks to predict multi-shell data from single-shell data and used the predicted data to estimate the NODDI model [114]. They achieved satisfactory results on healthy brains as well as brain of MS patients. In a similar work, Chen et al. [95] successfully trained an multilayer perceptron (MLP) to predict six-shell data from two-shell data. A closely related application is imputation of missing data, which is of particular importance because some dMRI measurements often have to be discarded due to strong noise, artifacts, or excessive motion, such as in young or noncooperative subjects. To address this problem, Hong et al. developed a method based on graph CNNs [116]. They represented the relation between the measurements in the voxel space and q-space using a graph, and trained a residual graph CNN to learn this relation. They trained the model using adversarial techniques and successfully applied it on data from healthy infants between 0 and 12 months old. This method was able to reconstruct accurate maps of generalized FA (GFA) and FOD from five-fold-reduced slice acquisitions.
IV-C Data harmonization
Data harmonization is crucial for reliable analysis and comparison of inter-scanner and inter-site dMRI data. It has become an increasingly more relevant problem as larger and geographically/demographically more diverse datasets are used in multi-center neuroimaging studies. Statistical methods such as ComBat [117] and RISH [118] have achieved considerable success in addressing this problem. However, they have important limitations. For example, they depend on data from matched subjects in the reference and target datasets, and their performance deteriorates when the differences between reference and target datasets are nonlinear [119].
TABLE I: A listing of some of the machine learning-based methods for data harmonization in dMRI.
  Method Summary of methodology and results
  Semi-supervised contrastive learning using CNNs [120] This method uses paired data to learn subject-specific and acquisition-specific representations. A decoder is trained to map the subject-specific representations to the target contrast. The method is shown to be superior to a method based on CycleGAN [121] as well as interpolation based on SHORE [122]. It can handle heterogeneous data from multiple sites, different acquisitions protocols, and demographics.
Null space learning with an MLP architecture [123] The model is trained using paired dMRI scans of human subjects as well as dMRI-histology data from squirrel monkeys. The loss function is designed to encourage accurate FOD estimation and consistent scan-rescan GFA. Experiments show that this method has high accuracy in FOD estimation and high reproducibility on test data from unseen scanners. Experiments also show that this approach results in higher FOD estimation accuracy across 1.5T and 3T scanners.
Variational auto-encoder [20] The method is based on learning a latent representation that is invariant to scanner, scanning protocol, or similar confounders. The latent representation can then be used to reconstruct the image content that is stripped of those factors. The method can be trained without paired scans. Training involves an adversarial loss that attempts to predict the source (scanner, etc.) of the acquisition. Results show that, compared with a well-established conventional technique [124], this new method achieves superior results in terms of several parameters including FA, MD, and fiber orientation.
Residual CNN [125] This method is based on predicting the spherical harmonics representation of the dMRI signal in the target domain from that of the source domain. It works on 3D patches of size 3. A final non-learned projection in the spherical harmonics space is needed to make sure fiber orientations are not changed due to the intensity harmonization. This method requires that the same subjects be scanned in the source and target scanner/acquisition protocol. Evaluations have shown that this method can achieve effective harmonization of the dMRI signal, FA, and MD.
Hierarchical CNNs [126] The focus of this work is on harmonization of DKI measures. The method requires a set of subjects to be scanned with both target and source scanners/acquisition protocols. It computes the DKI metrics using the data in the target domain. The network is trained to map the dMRI data in the source domain directly to the DKI measures in the target domain, after nonlinear intra-subject registration. This method reduced the inter-scanner variation in DKI measures by 51-66%.
CNN with a scan-rescan consistency loss [127] This method is proposed and evaluated specifically for FOD estimation, although it seems to be directly applicable for estimating any other parameter. A consistency loss penalizes the model for divergent FOD predictions from different scans of the same subject. It achieves better FOD estimation accuracy and inter-scanner consistency than standard techniques on external test data.
Ensemble of different neural networks [128] The authors present their method as a multi-task learning approach. Essentially, a set of neural networks, which can have different architectures and trained for different prediction tasks, are combined via training a set of additional neural networks that utilize the features learned by these networks to predict the parameter of interest. Compared with state of the art deep learning techniques, this method achieved better dMRI signal prediction.
Adaptive dictionary learning [129] A dictionary is learned on the set of reference datasets/scanner(s). It is assumed that representing a test scanner’s data in this dictionary automatically harmonizes the data towards the reference dictionary by suppressing the features that are specific to the test data. The method does not require paired subjects in the source and reference datasets. Successful harmonization results are reported in terms of FA, MD, and rotationally invariant spherical harmonics representations.
Recently, various machine learning-based methods have been proposed, but they still have not achieved the popularity of methods such as ComBat and RISH. A recent study compared a range of interpolation methods, statistical regression techniques, and deep learning methods for cross-scanner and cross-protocol multi-shell dMRI data harmonization [41]. A regression method based on [130] performed better than all deep learning methods. However, some of the deep learning methods were among the best performing techniques. The authors of [41] hypothesized that the performance of the deep learning methods may significantly improve with larger training datasets than the 10 training subjects used in that work. Tax et al. evaluated five different learning-based harmonization methods (including four neural networks and a dictionary-based technique) on data from different scanners and with different maximum gradient strengths [131]. Their results showed that overall these learning-based methods were successful in harmonizing cross-scanner and cross-protocol data, although no comparison with the state of the art non-learning methods was performed. Their analysis also showed that the learning-based methods were more effective on isotropic measures such as MD than on anisotropic measures such as FA. They attribute this behavior to the possibility that, because the spatial variations in isotropic measures are less abrupt, imperfect spatial alignments may be less harmful when applying machine learning methods on imperfectly-registered pairs of dMRI volumes.
Table I lists some of the existing machine learning methods for dMRI data harmonization. In general, these methods offer higher flexibility in modeling the sources of heterogeneity and in handling data from unpaired subjects. For example, one study developed a method based on representing the data in a disentangled latent space that allowed for separating the effects of anatomy and acquisition [120]. Similarly, variational auto-encoders have been proposed to harmonize dMRI data by learning a latent representation that is invariant to site/protocol-specific factors [20]. Focusing on the specific task of FOD estimation, another study developed a method, named null space deep network, that seamlessly integrated a small dMRI dataset with histology ground truth and a large scan-rescan dMRI dataset without histology ground truth [123]. The method showed superior FOD estimation accuracy on data from scanners that had not been included in the training set. A similar method was applied to harmonize structural connectivity metrics on a two-center dataset [132]. The authors found that optimal harmonization was achieved when the deep learning method was applied to harmonize the dMRI data at a voxel level followed by applying ComBat [117] to harmonize the structural connectivity metrics. Another work from the same research team, again focusing on FOD prediction, proposed an extra loss term to encourage consistency of FOD estimations computed from scan-rescan data of the same subject [133]. This new loss function improved the generalizability of FOD estimation as well as downstream tasks such as structural connectivity assessment. Blumberg et al. have proposed a multi-task learning strategy, where a neural network is trained to combine the information learned by an ensemble of different neural networks [128]. The individual neural networks in the ensemble may have been trained on separate datasets, possibly for completely different tasks. The added neural network model uses the high-level features learned by the ensemble to perform data harmonization. The authors argue that this approach can be especially effective in scenarios where one or some of the datasets are very small.
IV-D Microstructure mapping
The brain white matter consists of a network of neuronal fibers that are supported by other cellular components [134, 32]. Microstructural characteristics of the white matter tissue such as myelination influence its physical properties such as viscosity, density, and permeability [135, 136]. These physical properties, in turn, can influence the diffusion of water molecules. Although the size of these microstructural elements are in the micrometer range, dMRI can probe the tissue microstructure because dMRI signal is sensitive to the displacement of water molecules at the micrometer scale [2, 137]. Therefore, the dMRI signal can be an indicator of changes/variations in brain tissue microstructure due to normal/abnormal development or diseases [138, 139, 140, 5, 6].
@@ IV-D Microstructure mapping
There have been many efforts and much progress in developing bio-physical models that relate the measured dMRI signal to the microstructure of brain tissue [32, 31]. Advanced models rely on specialized measurement protocols and complex numerical optimization methods. The underlying estimation problem is typically nonlinear, sensitive to measurement noise and initialization, and may be unstable and computationally intensive [141, 142]. Accuracy and precision of model fitting is hard to verify and can significantly depend on the optimization algorithm used to compute the fitting [141]. To avoid local minima, some works have resorted to computationally intensive approaches such as grid search, multi-start methods, cascade optimization, and stochastic optimization techniques [143, 144, 142, 141]. There is a great interest in developing methods, such as those based on dictionary matching [145, 146], to reduce the computational time without compromising the estimation accuracy. A good example of such methods is AMICO [146], which reformulates the microstructure estimation equations as convex optimization problems that can be solved much faster.
IV-D1 Potential advantages of machine learning methods
In recent years, machine learning has increasingly been applied to these estimation tasks. Overall, five main justifications have been cited for preferring a machine learning-based approach for this application.
  1. 1.
@@ IV-D1 Potential advantages of machine learning methods
    Unlike the conventional methods that presume a known fixed relationship between the dMRI measurements and the target parameter, machine learning methods can learn this relationship from data [42, 16]. For certain microstructural parameters, such as residence time of water inside axons, it is believed that existing mathematical models that express the relationship with the dMRI signal are either too simplistic or intractable, and that the existing numerical forward models are computationally too expensive to be used in estimating the parameters from data [60, 45]. Fick et al. show that, for axonal diameter estimation, existing signal models fail outside a narrow range of diameters while a machine learning method can achieve accurate estimation for the whole range of diameters in the data [42].
  2. 2.
    Conventional methods often involve several steps with no feedback from the later steps to the earlier steps [16, 141]. As a result, these methods are hard to design and optimize. Machine learning methods such as deep neural networks, on the other hand, may be optimized end-to-end as a single processing step.
  3. 3.
    Machine learning methods can be much faster than numerical optimization routines. Again, such is the case for neural networks that run on GPUs [52, 53, 21].
  4. 4.
    Standard methods perform the model fitting in a voxel-wise manner, which fails to exploit the spatial correlations to improve the estimation accuracy. Machine learning models can effectively learn complex spatial correlations directly from data and leverage this knowledge to improve the estimation accuracy [147].
  5. 5.
    With machine learning methods it is typically much easier to incorporate prior knowledge or additional information such as constraints on the parameters to be estimated or other MRI contrasts.
Tables II, III, and IV list some of the recent works that have employed machine learning to estimate, respectively, DTI, DKI, and NODDI parameters. A summary of the methods developed in these works and their experimental results is presented below.
@@
TABLE II: Machine learning methods for estimating the diffusion tensor or its derived parameters such as FA and MD.
  Method Summary of methodology and results
@@
  SuperDTI [55] This work reports accurate estimation of FA, MD, and the main diffusion tensor eigenvector from six measurements. The model trained on healthy HCP brains works well on pathological brains. Evaluations included qualitative and quantitative tractography assessment.
DeepDTI [46] This method trains a CNN to compute dMRI data residuals. Specifically, it computes the residuals between under-sampled ($n=6$) data and high-quality targets computed from densely-sampled data. The diffusion tensor is then computed with a standard method. Compared with conventional estimation techniques, this method reduced the number of measurements by a factor of 3.3-4.6. It improved DTI estimation, DTI-based tractography, and tract-specific analysis on twenty prominent tracts. Only HCP data is used for validation.
Aliotta et al. [148] This work reported higher FA and MD reconstruction accuracy and precision than standard estimation methods from as few as 3 diffusion-weighted measurements. With a model trained on data from 10 healthy subjects, FA-based delineation of brain tumors was more accurate than with a standard method.
DIFFnet [149] The main contribution of this work is a method to handle the measurements acquired with different schemes. The method simply projects and bins the the q-space data in standard orthogonal planes. Compared with standard methods, they report faster computation and lower error.
@@
Fetal DTI estimation [22] This work has reported reconstruction of fetal DTI with unprecedented accuracy. The method was trained using synthetic data generated with a novel pipeline that used both fetal in-utero data and scans of premature neonates. Evaluations included quantitative comparisons with conventional methods as well as detailed assessment by human experts.
@@
Patch-CNN [150] The proposed method, a patch-wise CNN, reduces the required number of measurements by a factor of two compared with standard estimation methods. The estimated DTI maps were used to trace major white matter tracts with high accuracy.
@@
Transformer-based DTI estimation [51] This work used transformer networks to learn the spatial correlations in dMRI signal and in diffusion tensor. The method reconstructed the diffusion tensor with superior accuracy while reducing the required number of measurements by factors of 5-15. Evaluations include tractography and structural connectivity.
Cardiac DTI estimation (FG-Net) [151] This study has achieved accurate cardiac DTI estimation from six dMRI measurements. It includes a basic FCN that estimates the dMRI data for additional directions than the six measured. The method predicts DTI metrics more accurately than conventional methods on ex-vivo data.
Cardiac DTI [50] The model (a residual CNN) trained with purely synthetic data performs well on synthetic as well as in-vivo test data for cardiac DTI estimation. It is less prone to predicting implausible values and enables more accurate detection of tissue lesion.
SwinMR (cardiac DTI) [152] Deep learning methods can reconstruct cardiac DTI with k-space under-sampling rates of up to 4 without any significant quality reduction compared to reference. A transformer network is shown to achieve superior results than a CNN.
IV-D2 Method design